Write the slope-intercept equation for the line that passes through (-4, 6) and (2, -6). Please show all of your work
Answer by rfer(16322)

m=(-6-6)/(2--4)
m=-12/6
m=-2
----------------
y=mx+b
-6=-2*2+b
-6=-4+b
-2=b
--------------
y=-2x-2